OpenStudy (anonymous):

A rectangular fields perimeter is 92 yards with one side that measures 17 yards. The sides of the square field with the same area as the rectangular field measures how many yards?

OpenStudy (anonymous):

First we would need to find the other side of the rectangle. This is how we'll find the area of the rectangle. \[17+17=34\]\[92-34=58\]\[58\div2=29\]\[17\times29=493\]So, the other side of the rectangle is 29, so the area is 493. Do you follow so far?

OpenStudy (anonymous):

I'll continue to help after a response from you. When you're ready, just tag me.

OpenStudy (anonymous):

Yes i get that part

OpenStudy (anonymous):

Ok, so since it's a square that we need to find the length of the side for we can just find the square root of 493. So, \[\sqrt{493}=?\]

OpenStudy (anonymous):

22.20

OpenStudy (anonymous):

Yes, but there is no reason for the extra 0 on the end. 22.20 is the same as 22.2

OpenStudy (anonymous):

oh ok. that makes sense

OpenStudy (anonymous):

how do you give someone a medal?